OpenStudy (anonymous):

how do I set up this word problem? Kay spends 250 min/wk exercising. Her ratio of time spent on aerobics to time spent on weight training is 3 to 2. How many minutes does she spend per week on each?

OpenStudy (regan):

3/2=250/x solve

OpenStudy (anonymous):

do I multiply or divide?

OpenStudy (anonymous):

cross muiltiply that :]

OpenStudy (anonymous):

3x + 2x = 250 5x = 250 x = 50 so 3x = 150 minutes on aerobics and 2x = 100 minutes on weight training

OpenStudy (anonymous):

It is not 3/2 = 250/x that would be if it said she spent 250 minutes doing aerobics

OpenStudy (anonymous):

3/250= 2/x gives the same answer..

OpenStudy (regan):

o i see, i dont really consider doing wieghts excersise

OpenStudy (anonymous):

Thank you for your help.